# A High-Performance Training-Free Pipeline for Robust Random Telegraph Signal Characterization via Adaptive Wavelet-Based Denoising and Bayesian Digitization Methods

### v1

@@ Abstract
Random telegraph signal (RTS) analysis is increasingly important for characterizing meaningful temporal fluctuations in physical, chemical, and biological systems. The simplest RTS arises from discrete stochastic switching events between two binary states, quantified by their transition amplitude and dwell times in each state. Quantitative analysis of RTSs provides valuable insights into microscopic processes such as charge trapping in semiconductors. However, analyzing RTS becomes considerably complex when signals exhibit multi-level structures or are corrupted by background white or pink noise. To address these challenges and support high-throughput RTS analysis, we introduce a modular and scalable signal processing pipeline combining dual-tree complex wavelet transform (DTCWT) denoising with a Bayesian digitization strategy. The adaptive DTCWT-based denoiser incorporates autonomous parameter selection rules for its decomposition level and thresholds, optimizing white noise suppression without manual tuning. Complementing this denoiser, our probabilistic digitizer effectively resolves binary trap states even under residual notorious background pink noise. The overall approach enables robust performance across varying noise levels and multi-trap scenarios, improving mean dwell time estimation and RTS reconstruction over classical and neural baselines. The method is up to 83× faster, training-free, and suitable for real-time or large-scale analysis. Evaluations confirm its generalizability, speed, and reliability, providing a strong foundation for future fully adaptive and automated RTS pipelines.
@@ I Introduction
As RTS signals grow more complex and involve multiple levels, sophisticated analysis protocols are in demand beyond generic signal processing methods, with a particular emphasis on denoising and digitization. Traditional denoising techniques include moving average (MA) filtering [13], fast Fourier transform (FFT) with frequency filtering, empirical mode decomposition [14], and various wavelet-based methods [15]. These techniques provide baseline denoising capabilities, especially effective for signals corrupted by white noise. Probabilistic models, such as hidden Markov models (HMMs) [16, 17, 18] and Gaussian mixture models (GMMs) [19], have also been applied, offering strong interpretability of RTS and quantification of RTS parameters. Unfortunately, these approaches often require careful parameter initialization, limiting their scalability and robustness in complex RTS environments. When confronting pink noise and multi-trap scenarios, these classical approaches fall short.
@@
While denoising has seen significant innovation with these advanced models, RTS digitization, the process of mapping denoised signals to discrete levels and decomposing them into binary sequences for individual traps, has received comparatively less attention. Especially, the task becomes particularly challenging when signals involve more than 2 levels. Current practice primarily relies on simple thresholding methods [25] or proximity-based peak-matching heuristics [20, 19]. Some probabilistic-based methods have been explored in the context of RTS noise analysis [26, 27], offering a way to model and interpret fluctuating signals in a statistically grounded manner. Bayesian approaches, particularly suited for handling uncertain and noisy data, are valuable for extracting meaningful information from RTSs, where noise often obscures state transitions [26, 27].
@@
Finally, a unified framework for evaluating RTS analysis performance remains missing from the literature, as prior studies typically assess only a subset of metrics, without a comprehensive or standardized evaluation. To address all these limitations, we introduce a comprehensive RTS signal processing pipeline that integrates denoising and digitization within a single modular framework, along with a suite of quality and performance metrics, including various novel measures including trap count error ($N_{\text{trap}}$ error), to thoroughly evaluate the ability of each algorithm to characterize individual RTS traps.
@@
Here we offer a comprehensive and modular RTS signal processing pipeline designed for robust denoising, high-fidelity digitization, and statistical characterization. At its core are two primary algorithmic contributions: an adaptive dual-tree complex wavelet transform (DTCWT) denoiser and a Bayesian digitizer. Each contribution is selected to address limitations observed in prior methods, particularly under high-noise conditions and multi-trap scenarios. High-noise conditions, we refer to settings with a wide range of noise levels, including severe background fluctuations that can obscure step transitions—a point elaborated further in the Methods section.
@@ II Results
To benchmark the pipeline, we synthesize a dataset consisting of 1,800 RTS samples, each with length L=100, [^0] L=100{,}000 time steps. These samples span evenly distributed across ground truth trap counts $N_{\text{trap}}=\{1,2,3\}$, two background noise types (white noise (wn) and pink noise (pn)) and noise levels, $\eta_{\text{wn}},\eta_{\text{pn}}\in[1\%,30\%]$. Figure 1c–e illustrate 1-trap, 2-trap, and 3-trap RTSs ($a(t)$) injected with three levels of white noise $\eta_{\text{wn}}=10\%,20\%,30\%$ respectively. As both the number of traps and noise level increase, KDE peaksin the left sub panel become less distinguishable, and the time-lag plots increasingly overlap colonies of points. For a 3-trap RTS in Figure 1e, only 6 KDE peaks are identified, out of the 8 expected levels if three traps are independent and mutually exclusive. Figure 1f–h correspond to 1-trap, 2-trap, and 3-trap RTSs masked by the same pink noise level, $\eta_{\text{wn}}=10\%$ in all three cases. Higher trap counts under pink noise present greater denoising challenges: residual low-frequency components can create spurious levels in the KDE. In Figure 1g, for instance, a true 2-trap RTS may be misclassified as a 3-trap RTS due to multiple sub-peaks. Beyond the difficulty of colored-noise masking, a further major difficulty lies in efficient processing of very long RTS signals. In high-resolution experimental measurements, nanoscale devices can generate datasets containing millions or even billions of time steps. This scale places heavy demands on both memory and computation, making it essential to design denoising and digitization algorithms that are not only accurate but also adaptable and computationally efficient.
@@
We emphasize that our RTS datasets are generated under more challenging conditions than prior studies [20, 19], with a tighter definition of background noise strength $\eta_{\text{wn(pn)}}$ in Methods. Furthermore, we incorporate scaled trap amplitudes, where each additional trap has a reduced strength relative to the dominant trap, to simulate sub-dominant traps and introduce strong pink noise that can fully mask weaker transitions, thereby stressing both the denoising and digitization stages (details in Methods). To demonstrate the advantages of our approach, we define two primary benchmark categories: quality and performance. These benchmarks quantitatively assess both the proposed and baseline methods in a controlled and reproducible manner, ensuring that both accuracy and computational efficiency are considered.
@@ Discussion
Building on these benchmarking results, we now turn to a closer examination of the underlying mechanisms that drive the desirable performance of our proposed pipeline under challenging RTS conditions. First, the DTCWT denoiser exhibits the features of shift invariance, directional selectivity in the wavelet domain (i.e., distinguishing signal components by orientation in frequency rather than by time or amplitude), and strong noise suppression while preserving RTS step transitions. Unlike previous implementation requiring manual parameter tuning [28], our version includes an automatic parameter selection mechanism based on raw signal statistics, making it adaptive to diverse background noise regimes. For KDE, we employ a prominence-based peak filtering step that suppresses spurious peaks caused by noise, thereby improving the $N_{\text{trap}}$ estimation. This refinement is not the primary focus of our cross-method comparisons, since KDE is used consistently across all baselines, but it contributes to more reliable trap identification when it combines with the DTCWT denoiser. Second, the Bayesian digitizer replaces traditional heuristic-based digitization methods (e.g., thresholding or proximity rules) with a probabilistic model that estimates the most likely level transitions, particularly when signal levels overlap or traps are occluded. Unlike heuristic rules, the Bayesian framework incorporates prior knowledge about the system, which constrains inference toward physically plausible transitions even in ambiguous regions. Moreover, the probabilistic nature of the approach does not force a hard decision at every point but instead evaluates the full posterior distribution over possible states. This yields a principled measure of uncertainty, reduces susceptibility to noise, and naturally extends to multi-trap scenarios where interactions between levels complicate deterministic digitization. This allows for more reliable digitization and scales well in multi-trap settings. The resulting binary or multi-level sequences enable the best possible quantification of physical RTS parameters, $\bar{\tau}_{\text{high}}$, $\bar{\tau}_{\text{low}}$ and $\Delta_{\text{RTS}}$.
@@
Figure 8: Examples of processed real RTS from carbon nanotube film (a-c) with noisy RTS (grey), denoised RTS (blue), and digitized RTS (red) on middle subplot, cropped for better visualization; kernel density estimation (KDE) plot on left subplot; time-lag plot on right subplot for the entire RTS duration.
@@
Beyond quantitative evaluation, the proposed approach offers several key practical advantages. Most notably, it does not require training of machine learning models, enabling it to generalize across a wide variety of RTS conditions without overfitting to specific data distributions. In contrast, neural-network-based methods like DAEUBL, although being powerful under ideal conditions, are often less reliable on unseen real RTS data and demand substantial expertise to train, validate, and deploy [28]. Our method is inherently modular, allowing users to plug in different stages if needed, while remaining robust with minimal intervention. We also evaluated our method on experimental RTS signals collected from a carbon nanotube (CNT) device, visualized in Figure 8. The device under test consists of a 500 nm-wide, 65 nm-thick multi-walled CNT film acting as a conduction channel, with two terminals for voltage bias. The current through the channel is recorded at a temperature of 9 K. In Figure 8a, the time-lag plot of this sample shows mostly round-shaped intensities, indicative of white noise. The RTS signal is clean and clearly exhibits discrete transitions, making it well-characterized as a single-trap RTS. The denoiser performs smoothly and confidently classifies the step levels. The sample in Figure 8b exhibits stretched intensities in the time-lag plot, consistent with pink noise. Although it is noisier, it remains interpretable. The KDE reveals two lower peak levels near 4.258. Depending on tuning, one could retain one of these levels to model the signal as a 2-trap RTS with three levels, or discard both and interpret it as a simplified single-trap RTS. This flexibility demonstrates the value of manual refinement in edge cases, although the autonomous output is already reasonable. The last sample in Figure 8c also displays mostly white noise characteristics in the time-lag plot. Although the KDE suggests a potential high-level peak, the denoised signal offers limited support for this level, particularly from $t=0$ to $t=5000$. The time-lag density confirms that this upper peak occurs much less frequently (e.g., lower green point density vs. purple), justifying its exclusion. With only two transitions observed at that level, it is difficult to confidently label it as a valid RTS state. Nevertheless, the conservative decision by the denoiser here is acceptable for autonomous operation, and such borderline cases can be manually reviewed if necessary.
@@ Methods
To emulate realistic experimental conditions, controlled noise is added to the ground truth RTS signal, which can be either single-trap or multi-trap. $A(t)$ denotes the ground truth RTS amplitude at time or signal step $t$, to which noise is added with intensity scaled by a factor $\eta_{\text{wn}}$. Let $\Delta A=\max(A)-\min(A)$ be the RTS maximum amplitude range. Upon synthesizing the white noise to the ground truth signal, we define the white noise $W(t)\sim\mathcal{N}(0,\sigma_{W}^{2})$ where $\mathcal{N}(0,\sigma_{W}^{2})$ denotes a Gaussian distribution with zero mean and variance $\sigma_{W}^{2}$. The standard deviation is $std(W(t))=\sigma_{W}=\Delta A\cdot\eta_{\text{wn}}$. The synthesized RTS with white noise is then expressed as $a(t)=A(t)+W(t)$. To synthesize RTS samples with pink noise as a form of low-frequency noise, we replace the previously modeled Gaussian white noise with pink noise, characterized by a power spectral density (PSD) that follows a $1/f^{\alpha}$ trend, where $f$ denotes frequency and $\alpha$ is the power-law exponent. We assume the standard pink noise with $\alpha=1$. The pink noise $P(t)$, when the given noise level is $\eta_{\text{pn}}$, is derived by applying a power-law filter to white noise $W(t)$ where $\eta_{\text{wn}}=\eta_{\text{pn}}$. The filtered output $\text{PL}(t)$ has a $1/f$ PSD characteristic. Therefore, the pink noise term is defined as $P(t)=\sigma_{W}\cdot\text{PL}(t)$, where $\text{PL}(t)$ is a pink noise sample with $1/f$ spectral properties by first generating white noise. The resulting pink noise RTS thus becomes $a(t)=A(t)+P(t)$. The process of adding controlled noise can be visualized in Figure 1 from ground truth RTS (black) to noisy RTS (grey).
@@
Since the DTCWT denoising requires both a decomposition level and a threshold value, we conduct an exhaustive grid search over reasonable parameter ranges to identify optimal combinations. For each RTS signal, we duplicate the data across the full parameter grid and process each copy through the DTCWT denoising followed by digitization. We then compute the deviation between the extracted and ground-truth dwell times ($\bar{\tau}_{\text{low}}$, $\bar{\tau}_{\text{high}}$), selecting the parameter pair that minimized this error. From these experiments, we identify strong statistical correlations between the optimal DTCWT parameters and simple measurable signal properties. Specifically, we find that the decomposition level $K$ correlates with the signal length $L$, while the wavelet coefficient threshold $T$ is best predicted by the spectral entropy $H_{S}$ of the signal. This enables us to derive the following predictive rules using linear regression fits over the grid search results:
@@
Here, $H_{S}$ is defined as $H_{S}=-\sum_{k=1}^{N}P_{k}\log P_{k}$, where $N$ is the total number of frequency bins obtained from Welch’s method, $X_{k}$ is the power spectral density coefficient at the $k$-th frequency bin, and $P_{k}=|X_{k}|^{2}/\sum_{i=1}^{N}|X_{i}|^{2}$ corresponding normalized power spectral density of the signal. For clarity in this study, we refer to our fully automated DTCWT-based denoiser with adaptive selection of $K$ and $T$ simply as DTCWT, distinguishing it from the original transform and from basic denoising implementations that require explicit parameter specification.
@@
After denoising, the next step in the pipeline is to identify the stable amplitude levels present in the signal. For this purpose, we apply KDE, a non-parametric technique used to estimate the probability density function of the signal amplitudes. KDE operates by placing a smooth kernel, typically a Gaussian function, over each data point in the denoised signal, and summing the resulting distributions to form a continuous density function. This approach generates a smooth approximation of the amplitude distribution, which is particularly well-suited for noisy or fluctuation-heavy signals like RTS. Peak identification in the KDE result is performed using prominence-aware peak-finding algorithms. These methods assess not only the height of peaks but also their relative separation from neighboring fluctuations, providing robustness in noisy scenarios. When analyzing known 1-trap signals, it is possible to apply a forced mode to extract exactly two dominant peaks, ensuring that the digitization process focuses on the most physically meaningful states. However, this is not used throughout the entirety of this study to ensure no intervention in executing the protocols, meaning that it is possible to recognize a 1-trap RTS as a multi-trap when too much residual noise is present after denoising. Compared to traditional histograms, which can produce misleading visualizations due to arbitrary bin widths or boundaries, KDE offers a smoother and more faithful representation of the underlying distribution. This is especially useful when identifying discrete RTS levels, as noise and sampling artifacts can otherwise obscure the presence of multiple states. In the context of RTS analysis, KDE enables clearer visualization and quantification of distinct amplitude levels—corresponding to the discrete states of traps. These appear as peaks in the KDE output. For example, a 1-trap RTS typically produces two prominent peaks, corresponding to the high and low states. These peak positions are then used in the digitization step to assign each sample in the signal to its most probable discrete state. As illustrated in Figure 1c–h, the left subplots show the KDE results of the denoised RTS signals, where the density curves reveal the dominant peaks used to determine trap amplitude levels. The middle subplots then illustrate the progression from denoised RTS (blue) to digitized RTS (red). Unlike a histogram, which may display spurious gaps or bins that obscure these levels, the KDE plot offers a continuous and interpretable view that emphasizes the underlying state levels. This clarity is critical for downstream digitization and for accurate extraction of switching behavior.

### v2

@@ Abstract
[…]
@@ I Introduction
As RTS signals grow more complex and involve multiple levels, sophisticated analysis protocols are in demand beyond generic signal processing methods, with a particular emphasis on denoising and digitization. Traditional denoising techniques include moving average (MA) filtering [33], fast Fourier transform (FFT) with frequency filtering, empirical mode decomposition [22], and various wavelet-based methods [21]. These techniques provide baseline denoising capabilities, especially effective for signals corrupted by white noise. Wavelet-based approaches have also been widely employed in other signal denoising and multiscale analysis applications involving discrete and multilevel data, including large-scale N-body and cosmological simulations, where they have been employed to mitigate discreteness effects and analyze scale-dependent structure [28, 29, 27]. Probabilistic models, such as hidden Markov models (HMMs) [10, 8, 17] and Gaussian mixture models (GMMs) [26], have also been applied to RTS analysis, offering strong interpretability of RTS and quantification of RTS parameters. Unfortunately, these approaches often require careful parameter initialization, limiting their scalability and robustness in complex RTS environments. When confronting pink noise and multi-trap scenarios, these classical approaches fall short.
@@
While denoising has seen significant innovation with these advanced neural-based and wavelet-based models, RTS digitization, the process of mapping denoised signals to discrete levels and decomposing them into binary sequences for individual traps, has received comparatively less systematic attention. Especially, the task becomes particularly challenging when signals involve more than 2 levels. Current practice primarily relies on simple thresholding methods [18] or proximity-based peak-matching heuristics [9, 26]. Some probabilistic-based methods have been explored in the context of RTS noise analysis [2, 4], offering a way to model and interpret fluctuating signals in a statistically grounded manner. Bayesian approaches, particularly suited for handling uncertain and noisy data, are valuable for extracting meaningful information from RTSs, where noise often obscures state transitions [2, 4], but existing formulations typically rely on complex models or require extensive tuning and prior assumptions [25, 5].
@@
Finally, a unified framework for evaluating RTS analysis performance remains missing from the literature, as prior studies typically assess only a subset of metrics, without a comprehensive or standardized evaluation protocol. To address this gap, we introduce a fully specified RTS signal processing pipeline that integrates denoising and digitization within a single modular framework, together with a comprehensive suite of quality and performance metrics. In particular, we define a 7-category benchmarking regime including various novel measures such as trap count error ($N_{\text{trap}}$ error), to thoroughly evaluate the ability of each algorithm to characterize individual RTS traps in a systematic and reproducible manner.
@@
Here we offer a comprehensive training-free and modular RTS analysis pipeline designed for robust denoising, high-fidelity digitization, and statistical characterization. At its core are two primary algorithmic contributions: an adaptive dual-tree complex wavelet transform (DTCWT) denoiser and a Bayesian digitizer. Each contribution is selected to address limitations observed in prior methods, particularly under high-noise conditions and multi-trap scenarios. High-noise conditions, we refer to settings with a wide range of noise levels, including severe background fluctuations that can obscure step transitions—a point elaborated further in the Methods section. While wavelet-based denoising [24] and probabilistic inference [2, 4, 25, 5] have each been explored independently in prior RTS studies, our work adopts in its adaptive formulation and unified integration into an autonomous pipeline that operates without training data or learned models. Each RTS trace is processed independently, avoiding dependence on prior signal statistics and reducing the barrier for experimental adoption.
The proposed adaptive DTCWT denoiser leverages the approximate shift invariance and better transition preservation of the traditional DTCWT by incorporating automated decomposition-level selection and thresholding rules specific to RTS transitions. The benchmarking results show that these characteristics allow for efficient background noise suppression while preserving transition fidelity under a variety of noise conditions. A probabilistic model that interprets the denoised signal in terms of discrete amplitude levels found from KDE is then used to handle RTS digitization. The digitizer determines RTS levels using a Bayesian inference process that integrates data from the observed signal amplitude and prior state likelihoods, as opposed to heuristic thresholding or proximity-based peak matching. In addition to being computationally efficient and appropriate for large-scale RTS analysis, this method offers a lightweight and comprehensible substitute for more intricate generative models. Methods provides a detailed description of the formal probabilistic formulation and inference process.
@@ II Results
To benchmark the pipeline, we synthesize a dataset consisting of 1,800 RTS samples, each with length L=100, [^0] L=100{,}000 time steps. These samples span evenly distributed across ground truth trap counts $N_{\text{trap}}=\{1,2,3\}$, two background noise types (white noise (wn) and pink noise (pn)) and noise levels, $\eta_{\text{wn}},\eta_{\text{pn}}\in[1\%,30\%]$. Quantitative benchmarking is performed on synthetic RTS signals, as real experimental RTS data generally lack ground-truth annotations for trap number, transition amplitudes, and dwell times, making objective quantitative evaluation infeasible without extensive device-specific physical modeling. Figure 1c–e illustrate 1-trap, 2-trap, and 3-trap RTSs ($a(t)$) injected with three levels of white noise $\eta_{\text{wn}}=10\%,20\%,30\%$ respectively. As both the number of traps and noise level increase, KDE peaks in the left sub panel become less distinguishable, and the time-lag plots increasingly overlap colonies of points. For a 3-trap RTS in Figure 1e, only 6 KDE peaks are identified, out of the 8 expected levels if three traps are independent and mutually exclusive. Figure 1f–h correspond to 1-trap, 2-trap, and 3-trap RTSs masked by the same pink noise level, $\eta_{\text{wn}}=10\%$ in all three cases. Higher trap counts under pink noise present greater denoising challenges: residual low-frequency components can create spurious levels in the KDE. In Figure 1g, for instance, a true 2-trap RTS may be misclassified as a 3-trap RTS due to multiple sub-peaks. Synthetic RTS signals therefore play a critical role in enabling controlled and reproducible evaluation across noise types, trap counts, and signal complexities, allowing systematic stress testing of denoising and digitization performance that is not possible with real experimental data alone. Beyond the difficulty of colored-noise masking, a further major difficulty lies in the efficient processing of very long RTS signals. In high-resolution experimental measurements, nanoscale devices can generate datasets containing millions or even billions of time steps. This scale places heavy demands on both memory and computation, making it essential to design denoising and digitization algorithms that are not only accurate but also adaptable and computationally efficient.
@@
We emphasize that our RTS datasets are generated under more challenging conditions than prior studies [9, 26], with a tighter definition of background noise strength $\eta_{\text{wn(pn)}}$ in Methods. Furthermore, we incorporate scaled trap amplitudes, where each additional trap has a reduced strength relative to the dominant trap, to simulate sub-dominant traps and introduce strong pink noise that can fully mask weaker transitions, thereby stressing both the denoising and digitization stages (details in Methods). Real experimental RTS examples are included in this work to demonstrate practical applicability of the proposed pipeline rather than to provide quantitative validation, as meaningful numerical error metrics on real RTS data typically require expert annotation and detailed device-specific physical interpretation, often forming the focus of a standalone experimental study. To demonstrate the advantages of our approach, we define two primary benchmark categories: quality and performance. These benchmarks quantitatively assess both the proposed and baseline methods in a controlled and reproducible manner, ensuring that both accuracy and computational efficiency are considered.
@@ Discussion
Building on these benchmarking results, we now turn to a closer examination of the underlying mechanisms that drive the desirable performance of our proposed pipeline under challenging RTS conditions. First, the DTCWT denoiser exhibits the features of shift invariance, directional selectivity in the wavelet domain (i.e., distinguishing signal components by orientation in frequency rather than by time or amplitude), and strong noise suppression while preserving RTS step transitions. Unlike previous implementation requiring manual parameter tuning [3], our version includes an automatic parameter selection mechanism based on raw signal statistics, making it adaptive to diverse background noise regimes. For KDE, we employ a prominence-based peak filtering step that suppresses spurious peaks caused by noise, thereby improving the $N_{\text{trap}}$ estimation. This refinement is not the primary focus of our cross-method comparisons, since KDE is used consistently across all baselines, but it contributes to more reliable trap identification when it combines with the DTCWT denoiser. Second, the Bayesian digitizer replaces traditional heuristic-based digitization methods (e.g., thresholding or proximity rules) with a probabilistic latent-level model over KDE peaks. For each time step, we compute the full posterior distribution under a Gaussian emission model and then report the digitized label via maximum a posteriori (MAP) decoding. Along with temporal prior smoothing, this yields a principled measure of uncertainty, reduces susceptibility to noise, and naturally extends to multi-trap scenarios where interactions between levels complicate deterministic digitization. This allows for more reliable digitization and scales well in multi-trap settings. The resulting binary or multi-level sequences enable the best possible quantification of physical RTS parameters, $\bar{\tau}_{\text{high}}$, $\bar{\tau}_{\text{low}}$ and $\Delta_{\text{RTS}}$.
@@
Figure 8: Examples of processed real RTS from carbon nanotube film with noisy RTS (grey), denoised RTS (blue), and digitized RTS (red) on middle subplot, cropped for better visualization; kernel density estimation (KDE) plot on left subplot; time-lag plot on right subplot for the entire RTS duration. (a-c) RTS from three distinct samples (Samples 1–3) processed using fully autonomous DTCWT + Bayesian. (d) RTS from Sample 3 reprocessed using manually tuned DTCWT + Bayesian following expert recommendation on potential alternative trap behavior. (e) RTS from Sample 3 reprocessed using DAEUBL + Proximity.
@@
Beyond quantitative evaluation, the proposed approach offers several key practical advantages. Most notably, it does not require training of machine learning models, enabling it to generalize across a wide variety of RTS conditions without overfitting to specific data distributions. In contrast, neural-network-based methods like DAEUBL, although being powerful under ideal conditions, are often less reliable on unseen real RTS data and demand substantial expertise to train, validate, and deploy [3]. Our method is inherently modular, allowing users to plug in different stages if needed, while remaining robust with minimal intervention. We also evaluated our method on experimental RTS signals collected from a carbon nanotube (CNT) device, visualized in Figure 8. The device under test consists of a 500 nm-wide, 65 nm-thick multi-walled CNT film acting as a conduction channel, with two terminals for voltage bias. The current through the channel is recorded at a temperature of 9 K. We emphasize that these real-data examples are presented to demonstrate practical applicability rather than to provide a complete physical characterization of CNT trap dynamics, which typically requires device-specific studies such as gate-voltage or temperature dependence and is beyond the scope of this work. In Figure 8a, the time-lag plot of this sample shows mostly round-shaped intensities, indicative of white noise. The RTS signal is clean and clearly exhibits discrete transitions, making it well-characterized as a single-trap RTS. The denoiser performs smoothly and confidently classifies the step levels. The sample in Figure 8b exhibits stretched intensities in the time-lag plot, consistent with pink noise. Although it is noisier, it remains interpretable. The KDE reveals two lower peak levels near 4.258. Depending on tuning, one could retain one of these levels to model the signal as a 2-trap RTS with three levels, or discard both and interpret it as a simplified single-trap RTS. Such ambiguity reflects common experimental scenarios where expected trap behavior may depend on prior physical insight; the proposed pipeline provides a reasonable autonomous interpretation while allowing expert-guided refinement when desired. The last sample in Figure 8c also displays mostly white noise characteristics in the time-lag plot. Although the KDE suggests a potential high-level peak, the denoised signal offers limited support for this level, particularly from $t=0$ to $t=5000$. The time-lag density confirms that this upper peak occurs much less frequently, justifying its exclusion under fully autonomous operation. However, prior literature [26] has suggested that the fast switching behavior below approximately 5.20 may correspond to a separate trap; motivated by this, we include an additional tuned result in Figure 8d, where reduced denoising strength preserves these transitions and yields a multi-trap interpretation. This comparison illustrates a key strength of the proposed framework: it enables reliable autonomous analysis across diverse noise regimes, while remaining flexible enough to incorporate expert expectations about trap behavior when such knowledge is available. In contrast, DAEUBL shows reduced robustness on this unseen real RTS example Figure 8e, as its training and noise synthesis do not fully span the experimental conditions encountered here.
@@ Methods
Traps are assumed to be independent in order to provide a controlled baseline for benchmarking that isolates algorithmic performance without introducing device-specific trap interactions. While correlated traps may arise in certain experimental systems, the independent-trap assumption does not trivialize the analysis task. To emulate realistic experimental conditions, controlled noise is added to the ground truth RTS signal, which can be either single-trap or multi-trap. $A(t)$ denotes the ground truth RTS amplitude at time or signal step $t$, to which noise is added with intensity scaled by a factor $\eta_{\text{wn}}$. Let $\Delta A=\max(A)-\min(A)$ be the RTS maximum amplitude range. Upon synthesizing the white noise to the ground truth signal, we define the white noise $W(t)\sim\mathcal{N}(0,\sigma_{W}^{2})$ where $\mathcal{N}(0,\sigma_{W}^{2})$ denotes a Gaussian distribution with zero mean and variance $\sigma_{W}^{2}$. The standard deviation is $std(W(t))=\sigma_{W}=\Delta A\cdot\eta_{\text{wn}}$. The synthesized RTS with white noise is then expressed as $a(t)=A(t)+W(t)$. To synthesize RTS samples with pink noise as a form of low-frequency noise, we replace the previously modeled Gaussian white noise with pink noise, characterized by a power spectral density (PSD) that follows a $1/f^{\alpha}$ trend, where $f$ denotes frequency and $\alpha$ is the power-law exponent. We assume the standard pink noise with $\alpha=1$. The pink noise $P(t)$, when the given noise level is $\eta_{\text{pn}}$, is derived by applying a power-law filter to white noise $W(t)$ where $\eta_{\text{wn}}=\eta_{\text{pn}}$. The filtered output $\text{PL}(t)$ has a $1/f$ PSD characteristic. Therefore, the pink noise term is defined as $P(t)=\sigma_{W}\cdot\text{PL}(t)$, where $\text{PL}(t)$ is a pink noise sample with $1/f$ spectral properties by first generating white noise. The resulting pink noise RTS thus becomes $a(t)=A(t)+P(t)$. The process of adding controlled noise can be visualized in Figure 1 from ground truth RTS (black) to noisy RTS (grey). Although the noise is additive, the inclusion of strong low-frequency (pink) noise introduces substantial temporal correlations and background fluctuations that can fully mask weaker RTS transitions, reflecting a major challenge commonly encountered in experimental RTS measurements.
@@
Since the DTCWT denoising requires both a decomposition level and a threshold value, we conduct an exhaustive grid search over reasonable parameter ranges to identify optimal combinations. This procedure is performed on a small set of representative real RTS measurements for which reference estimates were available, and is not based on the synthetic benchmark data used for evaluation. For each RTS signal, the data are duplicated across the full parameter grid and processed through DTCWT denoising followed by digitization. We then compute the deviation between the extracted and reference mean dwell times ($\bar{\tau}_{\text{low}}$, $\bar{\tau}_{\text{high}}$), and identify parameter regions that jointly minimize errors in both states. From these experiments, we observe strong statistical correlations between the optimal DTCWT parameters and simple measurable properties of the raw RTS signal. Importantly, these relationships depend only on general signal statistics and not on semiconductor-specific operating conditions or device physics. Specifically, we find that the optimal decomposition level $K$ correlates with the signal length $L$, while the optimal wavelet coefficient threshold $T$ is best predicted by the spectral entropy $H_{S}$ of the signal. This enables us to derive the following predictive rules using linear regression fits over the grid search results:
@@
Here, $H_{S}$ is defined as $H_{S}=-\sum_{k=1}^{N}P_{k}\log P_{k}$, where $N$ is the total number of frequency bins obtained from Welch’s method, $X_{k}$ is the power spectral density coefficient at the $k$-th frequency bin, and $P_{k}=|X_{k}|^{2}/\sum_{i=1}^{N}|X_{i}|^{2}$ corresponding normalized power spectral density of the signal. Spectral entropy was found to provide the most consistent correlation with the optimal threshold among the signal statistics examined. The resulting rules enable automated selection of DTCWT denoising parameters directly from the raw signal, without requiring prior knowledge of the underlying semiconductor system. The automatic selection does not fail in the sense of halting the algorithm or producing undefined outputs; rather, under extreme noise conditions, limitations manifest as increased benchmark errors, reflecting fundamental RTS reconstruction limits shared by all denoising methods. For real experimental RTS signals, where no ground truth exists, these rules provide a principled and reproducible starting point that can be further refined by expert users if desired. For clarity in this study, we refer to our fully automated DTCWT-based denoiser with adaptive selection of $K$ and $T$ simply as DTCWT, distinguishing it from the original transform and from basic denoising implementations that require explicit parameter specification.
@@
After denoising, the next step in the pipeline is to identify the stable amplitude levels present in the signal. For this purpose, we apply KDE, a non-parametric technique used to estimate the probability density function of the signal amplitudes. KDE operates by placing a smooth kernel, typically a Gaussian function, over each data point in the denoised signal, and summing the resulting distributions to form a continuous density function. This approach generates a smooth approximation of the amplitude distribution, which is particularly well-suited for noisy or fluctuation-heavy signals like RTS. Because KDE peak detection is sensitive to bandwidth selection, we use a fixed rule-of-thumb bandwidth to ensure reproducibility across all benchmark signals. Initially, KDE is applied to one-dimensional RTS amplitude samples ($d=1$), with the signal length $L$ corresponding to the number of samples. While Scott’s rule would prescribe a bandwidth scaling factor of $L^{-1/(d+4)}=L^{-1/5}$, we deliberately replace this data-dependent factor with a fixed empirical bandwidth (bw =0.020=0.020) applied uniformly after denoising and amplitude normalization. We found that using a fixed bandwidth yields more stable and interpretable KDE peak structures than bandwidths that shrink with signal length $L$. The value 0.020 was selected empirically as a conservative smoothing scale that balances suppression of residual fluctuations with preservation of distinct RTS levels. Within a reasonable range around this value, the number, ordering, and prominence of dominant KDE peaks remained stable in our tests. By avoiding signal-length–dependent bandwidth shrinkage inherent to rule-of-thumb methods such as Scott’s rule, this choice ensures consistent peak resolution across signals of varying length $L$ and noise level $\eta_{\text{wn}}$ or $\eta_{\text{pn}}$ without per-signal tuning. Peak identification in the KDE result is performed using prominence-aware peak-finding algorithms. These methods assess not only the height of peaks but also their relative separation from neighboring fluctuations, providing robustness in noisy scenarios. When analyzing known 1-trap signals, it is possible to apply a forced mode to extract exactly two dominant peaks, ensuring that the digitization process focuses on the most physically meaningful states. However, this is not used throughout the entirety of this study to ensure no intervention in executing the protocols, meaning that it is possible to recognize a 1-trap RTS as a multi-trap when too much residual noise is present after denoising. Compared to traditional histograms, which can produce misleading visualizations due to arbitrary bin widths or boundaries, KDE offers a smoother and more faithful representation of the underlying distribution. This is especially useful when identifying discrete RTS levels, as noise and sampling artifacts can otherwise obscure the presence of multiple states. In the context of RTS analysis, KDE enables clearer visualization and quantification of distinct amplitude levels—corresponding to the discrete states of traps. These appear as peaks in the KDE output. For example, a 1-trap RTS typically produces two prominent peaks, corresponding to the high and low states. These peak positions are then used in the digitization step to assign each sample in the signal to its most probable discrete state. As illustrated in Figure 1c–h, the left subplots show the KDE results of the denoised RTS signals, where the density curves reveal the dominant peaks used to determine trap amplitude levels. The middle subplots then illustrate the progression from denoised RTS (blue) to digitized RTS (red). Unlike a histogram, which may display spurious gaps or bins that obscure these levels, the KDE plot offers a continuous and interpretable view that emphasizes the underlying state levels. This clarity is critical for downstream digitization and for accurate extraction of switching behavior. We emphasize that KDE peak detection is intended to provide an automated and reproducible starting point for RTS analysis, while KDE may not perfectly identify all RTS levels in particularly challenging cases, such as those with extreme noise. These limitations are consistent with fundamental signal reconstruction constraints. In practice, such difficult experimental RTS traces often require expert inspection or manual refinement, which remains standard in experimental workflows. Accordingly, achieving zero error is neither expected nor required for the benchmarking goals of this study, and the impact of KDE uncertainties is explicitly captured by the proposed benchmarking metrics.